OpenStudy (anonymous):

Select the expression below that is identical to \[\frac{ 1 }{ 2 }\sec ^{2}\alpha \cot \alpha\] A.) csc2 alpha B.)1/2 C.)cos 2 alpha D.) sin alpha/2 E.) cos alpha/2

OpenStudy (anonymous):

Use these formulas and than simplify the expression. \[ \sec \alpha = \frac{1}{\cos \alpha}\] \[ \cot \alpha = \frac{\cos \alpha}{\sin \alpha}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }(\frac{ 1 }{ \cos \alpha })(\frac{ \cos \alpha }{ \sin \alpha })\] is that right?

OpenStudy (anonymous):

except the 1/cos a should be squared...

OpenStudy (anonymous):

Yes that's right.

OpenStudy (anonymous):

Now you can simplify a little bit.

OpenStudy (anonymous):

\[(\frac{ 1 }{ 2\cos ^{2}\alpha } )(\frac{ \cos \alpha }{ \sin \alpha})\]

OpenStudy (anonymous):

You can cancel the \(\cos \alpha\).

OpenStudy (anonymous):

so now its \[(\frac{ 1 }{ 2 \cos \alpha })(\frac{ 1 }{ \sin \alpha })\]

OpenStudy (anonymous):

OK, now you can use the fact that \(\sin(2\alpha) = 2\cos \alpha\sin\alpha \) and simplify the denominator.

OpenStudy (anonymous):

oh oh okay so then we get 1/sin 2 alpha

OpenStudy (anonymous):

That's right. But it can be written as \(\csc 2\alpha\).

OpenStudy (anonymous):

Because \(\frac{1}{\sin \alpha} = \csc \alpha\).

OpenStudy (anonymous):

Ugh i'm looking at my list of identities and I just like dont see the ones i need they you point them out and i feel like facepalming I'm so sorry.

OpenStudy (anonymous):

which woul dmake it answer C.)

OpenStudy (anonymous):

No, the correct answer is A. It's \(\csc 2\alpha\) not \(\cos 2\alpha\)

OpenStudy (anonymous):

Annnnd you're correct. :) thank you oodles. I'll have to look at my list more closely. I really appreciate the patience and help. I'm not exactly the brightest crayon in the box.

OpenStudy (anonymous):

Hope it helped. These trigonometric identities can be somehow tricky to remember.

OpenStudy (anonymous):

Just a little bit. Doesn't help that I dont care a bloody lick about 'em. Just got accept to the college I want to go to so I've hit the wall.